OpenStudy (anonymous):

sqrt14q times 2 sqrt49 Please show all work...

OpenStudy (anonymous):

@Maisa101

OpenStudy (anonymous):

\[\sqrt{14q} \times 2\sqrt{49}\] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So...wht next?!?

OpenStudy (anonymous):

okay what can you solve for first?

OpenStudy (kenljw):

sqrt1(4q) times 2 sqrt49) I'd factor coefficient 2sqrt(49)sqtt(14)(sqrt(q) 2X7sqr(14)sqrt(q)= 1414^1/2 q^1/2 = 14^3/2 q^1/2

OpenStudy (anonymous):